Grade 12th passMechanics

Water drops fall at regular intervals from a roof . At an instant when a drop is about to leave the roof,the separation between three successive drops below the roof are in the ratio .

Aditya Pathak
8 Years agoGrade 12th pass

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
D i s tan c e   t r a v e l l e d   i n   n t h   s e c o n d S = u + 1 2 g ( 2 n - 1 ) s 1 = 0 + 1 2 g ( 2 × 1 - 1 ) = g / 2 s 2 = 0 + 1 2 g ( 2 × 2 - 1 ) = 3 g / 2 s 3 = 0 + 1 2 g ( 2 × 3 - 1 ) = 5 g / 2 R a t i o   s 1 : s 2 : s 3 = 1 : 3 : 5